You are conducting a study to see if the probability of catching the flu this year is significantly different from 0.39. You use a significance level of α=0.02 α=0.02 . H0:p=0.39 H0:p=0.39 H1:p≠0.39 H1:p≠0.39 You obtain a sample of size n=427 n=427 in which there are 190 successes. What is the test statistic for this sample? (Report answer accurate to three decimal places.) test statistic = What is the p-value for this sample? (Report answer accurate to four decimal places.) p-value =
Answer #1

Given that, n = 427 and x = 190

sample proportion = 190/427 = 0.4450

The null and alternative hypotheses are,

H0 : p = 0.39 versus H1 : p ≠ 0.39

Test statistic is,

Test statistic = Z = 2.330

p-value = 2 * P(Z > 2.330) = 2 * 0.0099 = 0.0198

=> p-value = 0.0198